A share of stock sells for $31 today. The beta of the stock is 1 and the expected return on the market is 10 percent. The stock is expected to pay a dividend of $0.6 in one year. If the risk-free rate is 5.3 percent, what should the share price be in one year?
Answer #1

required return= risk-free rate +Beta*(MArket rate- risk-free rate )

=5.3+1*(10-5.3)

=10%

Required return=(D1/Current price)+Growth rate

0.1=(0.6/31)+Growth rate

Hence Growth rate=0.1-(0.6/31)

=0.080645161

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence share price in one year=$31*(1+0.080645161)

which is equal to

=33.50